### v1

# On the Noisy Road to Open Quantum Dynamics: The Place of Stochastic Hamiltonians

Abstract
Stochastic evolution underpins several approaches to the dynamics of open quantum systems, such as random modulation of Hamiltonian parameters, the stochastic Schrödinger equation (SSE), and the stochastic Liouville equation (SLE). In a stochastic formulation, the open-system problem is reduced from a coupled system-environment dynamics to an effective system-only description, with dissipative evolution recovered by ensemble averaging. In this work, we aim at a self-contained and accessible presentation of these approaches to further elaborate on their common roots in essential concepts of stochastic calculus and to delineate the conditions under which they are equivalent. We also discuss how different formulations naturally lead to different numerical time-integration schemes, better suited for either classical simulation platforms, based on finite-difference approximations, or quantum algorithms, that employ random unitary maps. Our analysis supplies a unified perspective and actionable recipes for classical and quantum implementations of stochastic evolution in the simulation of open quantum systems.
stochastic Hamiltonian; open quantum systems; stochastic Schrödinger equation; quantum trajectories; noisy quantum algorithms
preprint: APS/123-QED
I Introduction
Quantum systems are never fully isolated. Molecules and nano-structures, both natural and artificial, are embedded in some medium, such as solid-state matrices or fluids. Even almost-isolated systems in high vacuum, such as the neutral atoms and trapped ions used in quantum computing architectures, interact with the control apparatus and external fields. The surroundings and all the secondary degrees of freedom are part of the so-called environment of the system. The unavoidable interactions with the environment influence the expected ideal behavior of quantum systems, inducing decoherence and dissipation [petruccione2002, Vacchini2024OpenTheory], as they cannot be considered closed systems any longer. An effective treatment of open quantum systems then requires shifting from the unitary evolution of a wavefunction to a density matrix description, whose evolution does not follow a universal equation of motion but instead depends on the approximations and modeling strategies that one employs.
Two complementary perspectives are commonly adopted: in the first one, the density matrix describing the system of interest results from the partial trace over the environmental degrees of freedom and evolves deterministically under a quantum master equation. This approach requires starting from a microscopically explicit model of the environment and its interactions with the system, followed by either numerically-intensive reduction schemes, such as the hierarchical equations of motion (HEOM) or the time-evolving density matrix using the orthogonal polynomials algorithm (TEDOPA) [Tanimura1989TimeBath, Tanimura1990NonperturbativeBath, Tanimura2020NumericallyHEOM, Werner2016PositiveSystems, Lacroix2024MPSDynamics.jl:Dynamics, Cygorek2022SimulationEnvironments], or clear-cut approximations on the interaction strength and timescale separation, as in the seminal derivation of the reduced dynamics proposed by Redfield [Redfield1957, Redfield1965TheProcesses].
In the second approach, the density matrix is understood as the statistical average over an ensemble of stochastic pure-state trajectories. The latter viewpoint naturally motivates the introduction of stochastic Hamiltonians, in which the environmental influence is represented by time-dependent random fluctuations of system parameters. Averaging over realizations recovers the mixed-state dynamics of an open system, although the environment degrees of freedom are never considered explicitly. This framework offers a physically intuitive and computationally versatile route to include environmental effects in the dynamics, and has proven particularly powerful in the interpretation of spectroscopic observables where noise-induced dephasing and spectral diffusion play a central role and are reflected in the spectral lineshape [Anderson1954, Kubo1954AAbsorption, Kubo1963StochasticEquations, Haken1973AnMotion, Sanda2008StochasticExcitons]. More recently, a wide range of sophisticated techniques have been developed to control and drive the quantum state for information processing. When the control field is described within a semiclassical framework, the inclusion of fluctuations in the driving protocol directly maps onto a stochastic Hamiltonian description. This perspective is especially relevant in the context of noisy intermediate-scale quantum (NISQ) devices [Preskill2018QuantumBeyond, Brandhofer2021SpecialThem, Bharti2022NoisyAlgorithms], where unavoidable imperfections and control noise manifest as stochastic perturbations of the system Hamiltonian, thereby shaping both dynamics and computational fidelity [DeKeijzer2025QubitNoise, Baratz2025Data-DrivenDecomposition, Cialdi2017All-opticalChannels, Cialdi2019ExperimentalDynamics].
In this work, we start from the dynamics induced by a stochastic Hamiltonian [Kubo1963StochasticEquations, Haken1973AnMotion, Hasegawa1980StochasticProcesses], as it provides a particularly direct and versatile route. In this framework, the effective Hamiltonian $H^{\mathrm{eff}}_{t}$ is composed of the bare system Hamiltonian $H$ plus a randomly fluctuating term $H_{t}^{\mathrm{fluct}}$, so that we can write
$\hat{H}^{\mathrm{eff}}_{t}=\hat{H}+\hat{H}_{t}^{\mathrm{fluct}}=\hat{H}+Z_{t}\hat{R}$ (1)
where $\hat{R}$ is a Hermitian operator in the system state that defines the interaction with the environment, and $(Z_{t})_{t\geq 0}$ is either a continuous stochastic process or a discontinuous noise that models the effect of the environment on the system. Since the specific form of the system Hamiltonian $H$ is not relevant to our discussion, we will consider it as a generic time-independent Hermitian operator, but our conclusions can be easily generalized to the time-dependent case.
Direct introduction of random fluctuations in the Hamiltonians is not the only way of introducing stochastic dynamics in the quantum evolution. The stochastic Schrödinger equation (SSE) is a more general formalism in which the pure state of an open quantum system evolves according to a stochastic differential equation driven by noise terms [Gisin1984QuantumProcesses, Barchielli1991MeasurementsMechanics, Gisin1992TheSystems, Carmichael1993An18]. These noise terms were interpreted in connection with specific measurement protocols, “quantum jumps” or continuous measurement back-action [petruccione2002, Jacobs2006AMeasurement, Barchielli2009Quantum782]. Ultimately, the SSE is constructed so that the ensemble average over trajectories reproduces the deterministic evolution dictated by a master equation (in Lindblad form or non-Markovian extensions thereof [Barchielli2009Quantum782, Diosi1997TheSystems, Breuer1999StochasticEquations, Semina2014StochasticCases, Li2021MarkovianEquations, Gasbarri2018StochasticMaps]).
Renewed interest in stochastic methods in quantum dynamics has recently emerged in diverse contexts, ranging from molecular modeling [Coccia2018ProbingSystems, DallOsto2024StochasticSystems] to quantum technologies [Huang2020QubitNoise, DeKeijzer2025QubitNoise]. In this paper, we aim for a solid and clear analysis of these methods by uncovering and elaborating on their common roots in the basic concepts of stochastic calculus. We aim at offering a unified perspective that bridges methodological developments using stochastic Hamiltonians across chemical physics and quantum information science.
The stochastic formulation of the open-system dynamics replaces an explicit system–environment evolution with an effective system-only description, drastically reducing complexity. It also offers an efficient platform for the simulation of many-body systems (spin glasses, molecules, nanostructures), where computing the density-matrix evolution is often prohibitive due to the quadratic scaling of the state dimensions, compared to the linear scaling of the wavefunction. However, the shift from a deterministic to a stochastic dynamics affects both the formal apparatus and the interpretive framework.
In what follows, we present a self-contained analysis of stochastic processes in open quantum dynamics. Our primary goal is to clarify the interrelations among stochastic Hamiltonians, the stochastic Schrödinger equation (SSE), and the stochastic Liouville equation (SLE), and to delineate the conditions under which they are equivalent. To this end, we will make use of the two main interpretations of stochastic calculus, named after Itō and Stratonovich, presented in Section II.2, and point out explicitly the difference between a stochastic process and a noise. We further show how each formulation naturally maps to distinct numerical workflows, amenable to efficient classical trajectory integration or to quantum-algorithmic implementations. Figure 1 illustrates graphically the results presented in Sections III, LABEL:sec:ito and LABEL:sec:rode. First, in Section III, we discuss the lower blue pathway. Equation 6 is interpreted in the Stratonovich formalism as a viable way to implement quantum simulations, leading directly to the composition of universal gates in a quantum computing architecture. In LABEL:sec:ito, we follow the upper green branch of the schematic map and adopt the Itō interpretation, showing that it is the correct framework to obtain valid quantum master equations (QME). We also discuss how to write stochastic Schrödinger equations (SSE), and clarify how a stochastic Hamiltonian model is a special case of the SSE subject to more stringent constraints. In addition, we show that the Itō form of the evolution equations is easily treated with classical efficient time-integration schemes. A particular class of equations that can be cast indifferently in both Itō and Stratonovich frameworks is briefly discussed in LABEL:sec:rode. Finally, in LABEL:sec:sle, we show how we can work directly with stochastic density matrices instead of wavefunctions, moving back to the framework of earlier seminal works that employed stochastic Liouville equations [Kubo1963StochasticEquations, Haken1973AnMotion]. This allows for the inclusion of various effects, such as different dissipators and measurement apparatuses.
[…]
Figure 1: A schematic map of the two mathematical frameworks and relative workflows for the stochastic Hamiltonian applied to a wavefunction dynamics, based on the Itō (upper green branch) and Stratonovich (lower blue branch) interpretations. The Stratonovich pathway connects stochastic Hamiltonians to direct simulations of the density matrix dynamics through random unitary gates, while the Itō pathway provides the correct formulation for analytically obtaining quantum master equations. The random ordinary differential equation (RODE) case appears as a special limit when dealing with continuous stochastic processes as Hamiltonian fluctuations.
II Physical context and mathematical background
We now introduce the physical settings and mathematical frameworks on which our discussion builds. First, we describe how stochastic Hamiltonians arise from system–environment interactions and how they are modeled effectively within the system’s Hilbert space. Then, we review the key elements of stochastic calculus, with emphasis on the Itō and Stratonovich interpretations of stochastic differentials, which will play a central role in the developments of the subsequent sections and provide the mathematical language for the formulations that follow, as well as the definitions of stochastic process and noise.
II.1 Modeling of the stochastic Hamiltonian
We start from a system-environment interaction Hamiltonian of the type
$\hat{H}_{\text{int}}=\left(c\hat{S}+c^{*}\hat{S}^{\dagger}\right)\hat{B}$ (2)
where $\hat{S}$ acts on the system and $\hat{B}$ is a Hermitian operator acting on the environment. To confine the description within the Hilbert space of the system only, the environment operator $\hat{B}$ must be replaced with a function $B_{t}$. This is the first step of numerous embedding schemes, where the environment is reflected in a possibly time-dependent potential. Notice that when such a potential has a stochastic character, the expected change of the purity of the system state, typical of open quantum systems, can be recovered by averaging over different trajectories [Kubo1966TheTheorem, Gaspard1999Non-MarkovianEquation]. In some cases, the stochastic field can be considered as a surrogate environment, originating from the coupling Hamiltonian and the state of the original microscopic environment [Szankowski2020NoiseDynamics, Szankowski2021MeasuringNoise, Gu2019WhenNoise]. A prominent example used in triplet exciton dynamics is the replacement of a wide-bandwidth bosonic bath in the high-temperature limit with Gaussian white noise, which acts as a genuine surrogate field [Haken1973AnMotion].
The interaction Hamiltonian, now acting on the system subspace only, becomes
$\hat{H}_{t}^{\mathrm{fluct}}=Z_{t}\hat{S}+Z_{t}^{*}\hat{S}^{\dagger}$ (3)
where
$Z_{t}=cB_{t}.$ (4)
For real interaction coefficients $c$, we obtain real processes $Z_{t}$, allowing us to identify the operator
$\hat{R}=\hat{S}+\hat{S}^{\dagger},$ (5)
and retrieve eq. 1.
The dynamics of the reduced system, which originates from the evolution of the system plus environment under the Hamiltonian $\hat{H}^{\prime}=\hat{H}+\hat{H}_{\text{int}}$, unravels in terms of trajectories. We can write the Schrödinger equation for a system with Hamiltonian given by eq. 1 as
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}{t}}{\psi_{t}^{\mathrm{trj}}}=-i\left(\hat{H}+Z_{t}\hat{R}\right){\psi_{t}^{\mathrm{trj}}},$ (6)
where the superscript acknowledges the random nature of the single evolution, a trajectory. Depending on the formal structure of $Z_{t}$ and the target implementation, different formalisms should be used to ensure that eq. 6 is well defined, as we show below. From now on, to avoid cluttering the notation, we drop the hat symbol on the operators where the nature of each term is clear from the context.
The reduced density matrix of the open system is obtained as the mean density matrix over all the possible trajectories, that is
$\rho_{t}=\mathbb{E}\left[\rho^{\mathrm{trj}}_{t}\right]=\mathbb{E}\left[|\psi_{t}^{\mathrm{trj}}\rangle\langle\psi_{t}^{\mathrm{trj}}|\right],$ (7)
where $\rho^{\mathrm{trj}}_{t}$ is the density matrix of the system along a trajectory. In practice, the average on the ensemble $\Omega$ of trajectories $\omega\in\Omega$ is approximated by a finite sum as
$\mathbb{E}[\rho^{\mathrm{trj}}_{t}]=\lim_{n\to\infty}\frac{1}{n}\sum^{n}_{\omega=1}\rho_{t}^{(\omega)}\ \simeq\frac{1}{N}\sum^{N}_{\omega=1}\rho_{t}^{(\omega)}\,.$ (8)
II.2 Stochastic calculus tools
Introducing stochastic potentials to model random contributions to a system’s dynamics naturally calls for the use of stochastic calculus tools. Within this section, we provide a brief review of essential concepts. We recall the differences between the approaches by Itō and Stratonovich to stochastic calculus as they will be central to appreciate the application-oriented analysis of the following sections.
We introduce the definitions of stochastic differential and stochastic integral, needed to formalize the definitions of stochastic process and noise, in the two main frameworks that we need in the following. For a comprehensive and accessible introduction to the topic, we refer the reader to Ref. [Evans2013AnEquations].
We define a stochastic differential equation (SDE) as the formal expression
$\mathop{}\!\mathrm{d}{X_{t}}=f(t,X_{t})\mathop{}\!\mathrm{d}{t}+g(t,X_{t})\mathop{}\!\mathrm{d}{W_{t}}$ (9)
where $\mathop{}\!\mathrm{d}{X_{t}}$ is the stochastic differential of the process $X_{t}$, $f(t,X_{t})$ and $g(t,X_{t})$ are either functionals of the process or simple functions of time, $W_{t}$ is a Brownian motion (Wiener process), and $\mathop{}\!\mathrm{d}{W_{t}}$ its differential with scaling $\mathop{}\!\mathrm{d}{W_{t}}\sim\sqrt{\mathop{}\!\mathrm{d}{t}}$. Indeed, this type of notation is needed because the time derivative of the Wiener process is not a well-defined function; hence, the same is true for the time derivative of the generic process $X_{t}$.
Eq. 9 is a shorthand notation for the integral equation
${X_{t}}=X_{0}+\int_{0}^{t}f(s,X_{s})\mathop{}\!\mathrm{d}{s}+\int_{0}^{t}g(s,X_{s})\mathop{}\!\mathrm{d}{W_{s}},$ (10)
implying that we need to find a way to define an integral of the form
$\int_{0}^{T}g(t,X_{t})\mathop{}\!\mathrm{d}{W_{t}},$ (11)
that cannot be understood as an ordinary integral.
Expressing the integral as its Riemann-sum approximation on a partition of $m$ elements of the time interval $[0,T]$, we write
$\int_{0}^{T}g(t,X_{t})\mathop{}\!\mathrm{d}{W_{t}}\approx\sum_{k=0}^{m-1}g(\tau_{k},X_{\tau_{k}})\left(W(t_{k+1})-W(t_{k})\right)$ (12)
where $\tau_{k}$ is a point that lies within the subinterval $t_{k},t_{k+1}]$, i.e.
$\tau_{k}=\lambda t_{k+1}+(1-\lambda)t_{k}\quad\lambda\in[0,1].$ (13)
The choice of the parameter $\lambda$ defines different integral interpretations, introducing different frameworks [Evans2013AnEquations, Ito1975StochasticCalculus].
Itō’s integration is identified by $\lambda\!=\!0$, which corresponds to setting $\tau_{k}\!=t_{k}$, the left-hand point of each subinterval. This choice means that we do not need previous knowledge of the function at future values, in agreement with the idea that the stochastic potential provides unpredictable random perturbations to the system. In this interpretation of the stochastic differentials, the product rule for the differentials is not the ordinary one, but Itō’s product rule. For two differentials of the form $\mathop{}\!\mathrm{d}X_{i}=F_{i}\mathop{}\!\mathrm{d}t+G_{i}\mathop{}\!\mathrm{d}W_{t}$, the differential of the product is
$\mathop{}\!\mathrm{d}(X_{1}X_{2})=X_{2}\mathop{}\!\mathrm{d}X_{1}+X_{1}\mathop{}\!\mathrm{d}X_{2}+G_{1}G_{2}\mathop{}\!\mathrm{d}t,$ (14)
where, compared to ordinary calculus, we find the additional term $G_{1}G_{2}\mathop{}\!\mathrm{d}t$ due to the scaling $(\mathop{}\!\mathrm{d}{W})^{2}\sim\mathop{}\!\mathrm{d}{t}$.
Alternatively, among the infinite other possible choices, setting $\lambda=\frac{1}{2}$ leads to the Stratonovich integral definition. In this definition, the integral is time-symmetric, with both the function and the Wiener differential “sticking out toward the future” and the functional $g(X_{t},t)$ being now anticipating, or non-adapted. This means that the information at any time $t$ depends on some future-time values, but it agrees with the notion of a mid-point rule. Thus, in the differential Stratonovich notation, commonly identified by a circle before the stochastic differential $\circ\mathop{}\!\mathrm{d}{W}$, we can write
$\circ\mathop{}\!\mathrm{d}(X_{1}X_{2})=X_{1}\circ\mathop{}\!\mathrm{d}{X_{2}}+X_{2}\circ\mathop{}\!\mathrm{d}{X_{1}}$ (15)
where the usefulness of restoring the basic rules of differential calculus justifies the tricky anticipatory nature of this interpretation. A fundamental point now is the definition of the $\circ$ symbol as the symmetric Q-multiplication [Ito1975StochasticCalculus] which, for two time-dependent random variables $X_{t}$ and $Y_{t}$, is
$Y_{t}\circ\mathop{}\!\mathrm{d}X_{t}:=Y_{t}\mathop{}\!\mathrm{d}{X_{t}}+\frac{1}{2}\mathop{}\!\mathrm{d}X_{t}\mathop{}\!\mathrm{d}{Y_{t}}$ (16)
with the differential on the r.h.s. now expressed in the Itō interpretation. Then, by combinations of eqs. 16 and 15 we can transform a Stratonovich SDE into an Itō SDE.
Hereforth, we define a stochastic process as the integrated form of any Itō stochastic differential equation, $X=\int\mathop{}\!\mathrm{d}{X_{t}}$, that is the solution of eq. 10. The term noise corresponds instead to the formal time derivative of the stochastic process in a distributional sense, $\dot{X_{t}}=\mathop{}\!\mathrm{d}{X_{t}}/\mathop{}\!\mathrm{d}{t}$. Therefore, in the SDE formulation we define noise-inducing the stochastic differential $\mathop{}\!\mathrm{d}{X_{t}}$ itself. This definition can be understood by formally dividing eq. 9 by the time differential $\mathop{}\!\mathrm{d}{t}$ and obtaining a Langevin-like form, where the presence of a noisy drive is evident in the equation of motion for the process.
III Unitary propagation under stochastic Hamiltonians
Looking for the solution of eq. 6, we expect to be able to write a formal propagator of the form
$\mathcal{U}_{t,t_{0}}=\mathcal{T}_{+}\exp\left\{-i\int_{t_{0}}^{t}H^{\mathrm{eff}}_{s}\mathop{}\!\mathrm{d}{s}\right\}$ (17)
where $\mathcal{T}_{+}$ is the time-ordering operator, leading to solutions of the wavefunction evolution in the form
$\psi_{t}=\mathcal{U}_{t,t_{0}}\psi_{0}$ (18)
with initial value $\psi_{0}$, meaning that we should be able to propagate each trajectory through this unitary.
As anticipated in the introduction, we want to write eq. 6 in a formally correct way. We start writing the random potential as a noise, for simplicity the common white noise $\xi_{t}$ of intensity $\sigma$, by setting
$Z_{t}=\xi_{t}=\sigma\frac{\mathop{}\!\mathrm{d}{W_{t}}}{\mathop{}\!\mathrm{d}{t}},$ (19)
where $\mathop{}\!\mathrm{d}{W_{t}}$ is a proper stochastic differential and the formal time derivative is a distribution, the white noise distribution. Substituting this definition into eq. 6 and multiplying both sides by $\mathop{}\!\mathrm{d}{t}$, we obtain a mathematically correct formulation of the equation:
$\mathop{}\!\mathrm{d}\psi_{t}=-i{H}\psi_{t}\mathop{}\!\mathrm{d}{t}-i\sigma{R}\psi_{t}\circ\mathop{}\!\mathrm{d}{W_{t}}.$ (20)
The choice of the Stratonovich interpretation of the differential is fundamental, as it allows us to write the analytical solution to eq. 20, since the rules of ordinary calculus are valid, in the same fashion as for the solution of the Schrödinger equation
$\psi_{t}=\mathcal{T}_{+}\exp\left\{-i\int_{0}^{t}H\mathop{}\!\mathrm{d}{t}-i\sigma\int_{0}^{t}R\circ\mathop{}\!\mathrm{d}{W_{t}}\right\}\psi_{0}\,$, (21)
where we recognize in the time-ordered exponential the propagator of the dynamics, meaning that each trajectory can be propagated through this unitary. These solutions, proposed using white noise for the sake of clarity, can be easily generalized for colored noise, paying attention to the correct integration of the deterministic and stochastic parts.
Care must be used in the description of the dynamics of the mean density matrix of the open system, whose equation cannot be obtained by sticking to this formalism. Indeed, while eq. 7 holds true, if we take the product differential $\mathop{}\!\mathrm{d}{(\psi_{t}\psi_{t}^{\dagger})}$ and we average,
$\mathbb{E}\left[\mathop{}\!\mathrm{d}{(\psi_{t}\psi_{t}^{\dagger})}\right]=-i\left[H,\psi_{t}\psi_{t}^{\dagger}\right]\mathop{}\!\mathrm{d}{t}-i\mathbb{E}\left(\left[R,\psi_{t}\psi_{t}^{\dagger}\right]\circ\mathop{}\!\mathrm{d}{W_{t}}\right)$ (22)
we do not directly obtain the expected form of the master equation, a Lindblad form in the case of white noise stochastic driving. We refer to LABEL:sec:ito for the computation of the QME, for which one needs to work with Itō calculus.
III.1 Use of unitary propagators for implementation of quantum algorithms
In the framework discussed above, we obtain the expression of the propagator as a unitary operator, and norm-preservation is thus ensured trajectory-wise. We can obtain the numerical solution for the evolution in the time interval $[0,T]$, discretizing time into $S$ time steps $\tau=T/S$ and evolving the initial state through the first-order Magnus expansion [Magnus1954OnOperator, Blanes2009TheApplications] of the time-ordered propagator
$\mathcal{U}_{t}\approx\ \prod_{s=1}^{S}\ \mathcal{U}_{s\tau,(s-1)\tau}$ (23)
where each time-step unitary propagator is
$\begin{split}\mathcal{U}_{s\tau,(s-1)\tau}=\exp\Bigg\{&-i\int_{(s-1)\tau}^{s\tau}H\mathop{}\!\mathrm{d}{r}\\ &-i\sigma\int_{(s-1)\tau}^{s\tau}R\circ\mathop{}\!\mathrm{d}{W_{r}}\Bigg\}\end{split}$ (24)
and allows for numerical integration of every trajectory solution. For a time-independent system Hamiltonian and noise operator, the short-time evolution operators read
$\mathcal{U}_{\tau}=\exp\left\{-iH\tau-i\sigma R\Delta{W_{\tau}}\right\},$ (25)
where the white noise increment $\Delta{W_{\tau}}$ is drawn from a normal distribution of zero mean and standard deviation the square root of the time interval:
$\Delta{W\tau}\sim\mathcal{N}(0,\sqrt{\tau}).$ (26)
This numerical scheme, also used for implementations on classical computers [Jansen2006, Liang2012], allows the construction of quantum algorithms in digital quantum computer architectures without the overhead of auxiliary qubits. A simple example of a quantum algorithm with classical noise [gallina2022, Gallina2024FromDynamics] is depicted in LABEL:fig:circuit_QCNA, for a qubit evolution with Hamiltonian $H=\varepsilon\sigma_{z}+\Omega\sigma_{x}$, implemented using its first-order Suzuki-Trotter approximation [Trotter1959OnOperators, Suzuki1976GeneralizedProblems], and noise $\xi_{t}$ acting on the couplings through $R=\sigma_{x}$ (where $\sigma_{x}$ and $\sigma_{z}$ are the standard Pauli matrices).